# Predicting the Geolocation of Tweets Using transformer models on Customized Data

### v3

@@ 3.2 Model architecture
Table 3: Wrapper layer types by the number of outputs grouped by Spatial and Probabilistic output forms; Prediction Outcomes determines the number of predicted geographic points; Key Feature and Minor Features stand for the wrapper layers of the proposed models; Code column shows the corresponding model abbreviations.
Point - coordinate pair $\mathbf{\widehat{Y}}$; Weight - $\widehat{w}$ converted to the weights $W$; Cov - $\widehat{c}$ converted to the $\sigma_{\widehat{c}}$ parameter of the spherical covariance matrix $\mathbf{\Sigma}$
@@
Figure 2: Squared Euclidean Distance (SED) function surface on the axes of $\Delta Y_{lon}$ and $\Delta Y_{lat}$ as the error distances per longitude and latitude axes; upper horizontal gray surface indicates the empirical maximum of $L_{spat}$; red line indicates the strict minimum of 0 implied by the nature of Eq. (SED)
Figure 3: Negative Log-LikeliHood (NLLH) function surface on the axes of $D^{2}$ as the error distance and $\sigma_{\widehat{c}}$ as the uncertainty in $\widehat{\mu}$ of the Gaussian; red lines and gray surfaces indicate the reduction of $L_{prob}$ domain as a result of Eq. (LBSP) application
@@
To account for the variations in the output variables, each model type required an individual method for computing its loss. Such that models of the Geospatial type were utilizing solely $L_{spat}$, while models of the Probabilistic type were utilizing a combination of both $L_{spat}$ and $L_{prob}$. In terms of KF output evaluation, loss computation procedures for the SOP-type models are visualized in Figure 4, while computational graphs for the MOP-type models and multitask learning procedures can be found in Appendix B.
@@
Figure 4: Single Outcome Prediction (SOP) model loss functions computational graph including visualization of Squared Euclidean Distance (SED), Lower-Bounded SoftPlus (LBSP), and Negative Log-LikeliHood (NLLH) components
@@
To sum up, Section 3.2 covers a scope of all models reviewed in Table. To be more specific, it was the Probabilistic Multiple Outcomes Prediction (PMOP) model of 5 outcomes utilizing NON-GEO as the Key Feature and GEO-ONLY as the Minor Feature, which was finetuned using custom loss computation procedures described in Appendix B.
@@ 3.3 Results
In this section, the proposed models are evaluated by geospatial and probabilistic performance metrics on the worldwide and country-level granularity Twitter datasets. Importantly, the test evaluation was performed on the datasets filtered from bot users (posting more than 20 messages per day) and users utilized in the train datasets to ensure that the evaluation was unbiased and provided an accurate assessment of the models’ generalization capabilities.
The visualization of the post-processed prediction example for the best of the proposed models is shown below in Figure 5. The predicted set of weighted two-dimensional points and two-dimensional distribution with selected peaks were visualized for the geospatial and probabilistic models trained by the best approach. This example demonstrates a prediction of the randomly chosen text without a genuine truth location for the sake of comparison between two types of models trained on the Content and Context and producing an output of either multiple geolocation points (GMOP) or united distributions of location likelihood (PMOP). Considering the training data and procedures were the same for both models, there was a 15-degree difference between the GMOP and PMOP predictions. The indirect input without geographic references resulted in a relatively small shift of less than 2,000 km while maintaining the common guess of the US origins of the text.
Figure 5: Prediction examples of two models for the same text "CIA and FBI can track anyone, and you willingly give the data away"; both models were trained on the same worldwide dataset with key NON-GEO and minor GEO-ONLY text features, the number of prediction outcomes is 5;
above - GMOP significant (by weight) points as scatter plots; below - PMOP significant (by weight) Gaussian peaks as LLH and PDF plots
@@ 3.3.1 Worldwide evaluation metrics by model type
The models compared in Table 4 were trained on the local Twitter-World dataset composed of different Train Features (TF) and tested on both TEXT-ONLY and NON-GEO eValuation Features (VF). The most prevalent NG+GO training setup was using the hybrid approach with NON-GEO as the Key and GEO-ONLY as the Minor Features respectively. Metrics results for the MOP-type models were computed as the weighted linear combination of all outcomes.
Evaluation results of the worldwide models in Table 4 show the difference between geospatial and probabilistic models with a number of prediction outcomes ranging from 1 to 100. The test dataset consisted of 300K tweets from 143K users covering 63 languages and 226 countries. Note that both train and test datasets were the same for all models and contained tweets from both real and bot users.
@@
Table 4: Worldwide dataset (300,000 tweets) - results of models performance metrics; TF - Training Feature; VF - eValuation Feature;
NG+GO : NON-GEO + GEO-ONLY, A : ALL, NG : NON-GEO, TO : TEXT-ONLY;
all text features abbreviations are described in Table 1
u* - unlimited covariance output, $\sigma_{\widehat{c}}$ were obtained by the application of Eq. (SoftPlus) to the covariance parameter output $\widehat{c}$
Figure 6: Density (above) and Cumulative distribution (below) of the log error distance in km per outcome for the best model (Probabilistic, 5 outcomes, NON-GEO + GEO-ONLY) computed on the dataset of 100,000 worldwide tweets; median(above) and mean (below) errors and weights per outcome are noted in the legends below the graphs
@@
As for the Training Feature options, the last three models in Table 4 were trained on ALL ("text" + "user" + "place"), NON-GEO ("text" + "user"), and TEXT-ONLY ("text") features had higher SAE scores compared to the similar PMOP model of 5 outcomes trained on the proposed combination of NON-GEO and GEO-ONLY features, indicating underperformance.
Moreover, the increasing number of outcomes resulted in interchangeable spatial metrics scores, while revealing a modest growth of CAE and PRA metrics, as well as an undesirable decrease in the COV criteria. According to the analysis of the error distance per outcome depicted in Fig 6, up to 2 outcomes were assigned significant weights and only 73.76% predictions of the best outcome had median errors below 100 km.
@@ 3.3.2 Ablation study metrics
Table 5: Performance results of the model trained and evaluated on the Great Britain + France + Canada dataset using the hybrid approach (Content + Context) and PMOP-type output of 5 prediction outcomes for the estimation of the tweet location
Therefore, the ablation study on these specific countries effectively highlighted the model’s ability to handle ambiguity in language and demonstrated the influence of dataset composition on geolocation prediction accuracy. The ablation study results demonstrated that the model performed better on this limited dataset compared to the worldwide and American datasets. Notably, 96% of the predictions were within 161 km of the ground truth locations, which supports the suggestion of training country-scale models.
@@ 4 Discussion and Conclusion
This work was aimed at the examination of the machine learning techniques for solving the short text geolocation prediction task with NLP techniques employing a scope of BERT-based neural networks. The proposed framework demonstrates simplicity in terms of the trained model application as a tool for geotagging textual big data of short corpora. While the process of finetuning the models involved some complexity in loss functions, it is a one-time step that requires initial time and computing resources to develop the methodology of training the most accurate models.
@@
For instance, setup requirements for the proposed model application are the following: access to the local or remote model at https://huggingface.co/k4tel/geo-bert-multilingual; optionally dataset of text samples shorter than 512 tokens; and GPU or CPU node to launch the model.
@@
While the training routine requires access to the base model; dataset that contains labeled coordinates (lon, lat), text features for the arbitrary model input, and supplementary geotags of the locations; the usage of GPU is recommended due to the time and computing resources needed.
@@
One of the problems mentioned earlier was an alternative for the gazetteers - dictionaries of geographical indexes - used in the previous works containing accurate terms (location naming and LIWs) for geographical objects and their correct geographical coordinates. Considering the noisy user-generated data of the typical model inputs, the idea of feeding a consistent knowledge base of geographic objects into the model appeared practical and has been successfully implemented. The highest performance models were trained on multitask solving that offered an option of parallel learning with a total per-batch training loss as shown in Fig 7(d). The best strategy was split into key NON-GEO (always present "text" and "user") and minor GEO-ONLY ("place" present only in geo-tagged tweets) text features. In practice, every feature had its individual linear regression layer wrapping the base model that differed in the number of prediction outcomes, but matched in the per-feature loss function type (Geospatial/Probabilistic). The evaluation was performed using solely the key feature wrapper layer with no regard for the minor feature wrapper layer utilized only during the finetuning. This approach allowed for maintaining the format of the common model input and mapping the official place descriptions to the locations associated with the primary input texts. In terms of the described setup, the best performance results were achieved using the SOP-type minor loss and MOP-type key loss for 3-5 prediction outcomes.
@@
Figure 7: Loss logged during the training of the proposed models with single and multiple (5) outcomes and Geospatial and Probabilistic loss; KF - Key Feature, MF - Minor Feature; model graph lines by color:
orange - PMOP, blue - PMOP, red - GSOP, light blue - GMOP; model descriptions and abbreviations are provided in Table 3
Experiments have demonstrated that the geospatial loss of the minor feature GEO-ONLY shown in Fig 7(b) was declining much faster than the error distance of key feature NON-GEO shown in Fig 7(a) since the minor feature had less noisy and more persistent text data associated with the specific geolocations. The straightforward concatenation of the text content and metadata context of the "place" field would require undesirable randomization of all words in the string before tokenization such that models would not learn to only pay attention to the optional part ("place" metadata) of the input sequence. This statement is supported by the results in Table 4 which show that the model trained on ALL (union of "text", "user", and "place"), as the key feature has underperformed the model trained on the data split to key and minor features. Moreover, the model trained solely on key NON-GEO feature with no regard to the place context demonstrated even higher spatial errors during the evaluation. Thus the proposed methodology of multitask learning has demonstrated a significant improvement in the accuracy of the predictions and should be utilized if it’s possible.
Another proposed novelty was a lower-bound limitation of the covariance parameter $\sigma_{\widehat{c}}$ in the GMM output of the probabilistic models. Since the probabilistic loss function computed the negative log-likelihood of the labeled location to fit in the predicted Gaussian distributions, it depended on the Probability Density Function (PDF). In the case, $\sigma_{\widehat{c}}$ was approaching its minimum at 0, PDF exceeded 1 driving the loss function to the negative values as shown in Fig 3. In terms of probabilistic models, the total loss depended on both spatial and probabilistic components, therefore the latter gained a bigger momentum. As a result, the model paid less attention to the squared Euclidean error distance described in Eq. (SED) which negatively affected the geospatial accuracy. The suggested solution limited PDF to [0, 1] interval by setting the minimum of $\sigma_{\widehat{c}}$ to $\frac{1}{2\pi}$ such that Probabilistic loss shown in Fig 7(c) stays in the positive domain. Since the covariance parameter measures model uncertainty about the predicted location point, in the best-case scenarios, the probabilistic loss would depend mainly on the distance error as described in Section B.
@@
To address these privacy concerns, it is crucial for researchers, to anonymize or aggregate the data to minimize the risk of re-identification. The data utilized in this study has been protected with strict measures to prevent unauthorized access or breaches. Thus any collected Twitter dataset examples, containing metadata of user-profiles and geotags obtained from parsing the tweets, won’t be shared due to the potential privacy issues. By balancing the potential benefits of geolocation prediction with the protection of users’ privacy, it is possible to mitigate the privacy risks and promote responsible use of location data in Twitter research.
@@
In addition, there are several different world models that can be utilized for prediction output. While the Plate carrée projection is commonly used due to the Twitter geolocation stored in the WGS84 format, other projections such as the Robinson projection, conic projection, and Winkel-Tripel projection should be considered as possible alternatives in future studies.
In addition to BERT variations, the proposed machine learning techniques have the potential to be applied to other base models. In the realm of NLP, various state-of-the-art models have been developed to enhance text analysis tasks, particularly text classification that provides an output representation for the examined regression task. However, this would require adapting the implementation of the wrapper layer to accommodate the shape of the pooler output vector specific to the chosen model. Currently, the approach is designed for linear regression logic, transforming the vector of size 768, which is common to all BERT-based models, into a predefined number of continuous numerical values. Furthermore, in order to ensure suitability for the downstream task of sentence classification, it is important that the base models possess strong contextual understanding and the ability to capture semantic relationships within sentences.
@@
Overall, location-based sentiment analysis is an important tool for understanding public opinion, social dynamics and patterns, helping decision-makers and researchers to make data-driven decisions, and support the work of various sectors, from marketing to governance. This study provides the NLP-based approach for the estimation of geolocation by processing short text corpora such as social media posts on Twitter. The proposed solution utilizes multitask learning (key and minor features), context data (user and place metadata), and probabilistic output (GMM) to achieve higher spatial accuracy on the tasks of the tweet and the user’s home location prediction. Thus contributing to the field of big data analysis with a flexible geographical granularity setup of the custom BERT-based models.

### v4

@@ 3.2 Model architecture
Table 3: Wrapper layer types by the number of outputs grouped by Spatial (Spat) and Probabilistic (Prob) output forms; Prediction Outcomes determines the number of predicted geographic points; Key Feature and Minor Features stand for the wrapper layers of the proposed models;
Code column shows the abbreviations of different model configs.
Point refers to coordinate pair $\mathbf{\widehat{Y}}$; Weight is $\widehat{w}$ converted to the GMM weights $W$; and Cov corresponds to $\widehat{c}$ converted to the $\sigma_{\widehat{c}}$ parameter of the spherical covariance matrix $\mathbf{\Sigma}$
@@
Figure 2: Squared Euclidean Distance (SED) function surface on the axes of $\Delta Y_{lon}$ and $\Delta Y_{lat}$ as the error distances per longitude and latitude axes; upper horizontal gray surface indicates the empirical maximum of $L_{spat}$; red line indicates the strict minimum of 0 implied by the nature of Eq. (SED).
Figure 3: Negative Log-LikeliHood (NLLH) function surface on the axes of $D^{2}$ as the error distance and $\sigma_{\widehat{c}}$ as the uncertainty in $\widehat{\mu}$ of the Gaussian; red lines and gray surfaces indicate the reduction of $L_{prob}$ domain as a result of Eq. (LBSP) application.
@@
To account for the variations in the output variables, each model type required an individual method for computing its loss. Such that models of the Geospatial type were utilizing solely $L_{spat}$, while models of the Probabilistic type were utilizing a combination of both $L_{spat}$ and $L_{prob}$. In terms of KF output evaluation, loss computation procedures for the SOP-type models are visualized in Figure 4, while computational graphs for the MOP-type models and multitask learning procedures can be found in Appendix C.
@@
Figure 4: Single Outcome Prediction (SOP) model loss functions computational graph including visualization of Squared Euclidean Distance (SED), Lower-Bounded SoftPlus (LBSP), and Negative Log-LikeliHood (NLLH) components.
@@
To sum up, Section 3.2 covers a scope of all models reviewed in Table 3. According to the experimental results in Table 4, the most accurate model configuration was the Probabilistic Multiple Outcomes Prediction (PMOP) model of 5 outcomes utilizing NON-GEO as the Key Feature and GEO-ONLY as the Minor Feature, which was finetuned on a worldwide dataset of tweets using custom loss computation procedures described in Appendix C.
@@ 3.3 Results
In this section, the proposed models are evaluated by geospatial and probabilistic performance metrics in global (Table 4) and country-level (Table 5) granularity Twitter datasets. Importantly, the final evaluation was performed on the datasets filtered from bot users (posting more than 20 messages per day) and users utilized in the train datasets to ensure that the evaluation was unbiased and provided an accurate assessment of the models’ generalization capabilities.
The visualization of the post-processed prediction example for the best of the proposed models is shown below in Figure 5. The predicted set of weighted two-dimensional points and two-dimensional distribution with selected peaks were visualized for the Geospatial and Probabilistic models trained by the best approach. This example demonstrates a prediction of the randomly chosen text without a labeled location (genuine truth) for the sake of comparison between two types of models trained on the Content and Context and producing an output of either multiple geolocation points (GMOP) or united distributions of location likelihood (PMOP). Considering the training data and procedures were the same for both models, there was a 15-degree difference between the GMOP and PMOP predictions. The indirect input without geographic references resulted in a relatively small shift of less than 2,000 km while maintaining the common guess of the US origins of the text.
Figure 5: Prediction examples of two models, which were trained on the same worldwide dataset with key NON-GEO and minor GEO-ONLY text features, for the same text:
’CIA and FBI can track anyone, and you willingly give the data away’ with 5 outcomes sorted by significance (weight);
above: Geospatial GMOP points as scatter plots;
below: Probabilistic PMOP Gaussian peaks as LLH and PDF plots.
@@ 3.3.1 Worldwide evaluation metrics by model type
The models compared in Table 4 were trained on the local Twitter-World dataset composed of different Train Features (TF) and tested on both TEXT-ONLY and NON-GEO eValuation Features (VF). The most prevalent NG+GO training setup was using the hybrid approach with NON-GEO as the Key and GEO-ONLY as the Minor Features respectively. Metrics results for the MOP-type models were computed as the weighted linear combination of all outcomes, specific formulas for the MOP-type model can be found in Section B.
Evaluation results of the worldwide models in Table 4 show the difference between geospatial and probabilistic models with a number of prediction outcomes ranging from 1 to 100. The development dataset consisted of 300K tweets from 143K users covering 63 languages and 226 countries. Note that both train and development datasets were the same for all models and contained tweets from both real and bot users.
@@
Table 4: Worldwide dataset (300,000 tweets) results of models performance metrics;
TF stands for Training Feature, VF for eValuation Feature, and OUT defines number of outcomes while other model configuration parameters remain steady;
u*—unlimited covariance output, $\sigma_{\widehat{c}}$ were obtained by the application of Eq. (SoftPlus) to the covariance parameter output $\widehat{c}$
A: ALL or ’text’+’user’+’place’, NG: NON-GEO or ’text’+’user’, TO: TEXT-ONLY or ’text’,
NG+GO: NON-GEO + GEO-ONLY or (’text’+’user’)+’place’ where ’text’, ’user’, and ’place’ refer to the inputs defined in Table 1.
Figure 6: Density (above) and Cumulative distribution (below) of the log error distance in km per outcome for the best model (Probabilistic, 5 outcomes, NON-GEO + GEO-ONLY) computed on the dataset of 100,000 worldwide tweets; median (above) and mean (below) errors and weights per outcome are noted in the legends below the graphs.
@@
As for the Training Feature options, the last three models in Table 4 were trained on ALL (’text’ + ’user’ + ’place’), NON-GEO (’text’ + ’user’), and TEXT-ONLY (’text’) features had higher SAE scores compared to the similar PMOP model of 5 outcomes trained on the proposed combination of NON-GEO and GEO-ONLY features, indicating underperformance.
Moreover, the increasing number of outcomes resulted in interchangeable spatial metrics scores, while revealing a modest growth of CAE and PRA metrics, as well as an undesirable decrease in the COV criteria. According to the analysis of the error distance per outcome depicted in Figure 6, up to 2 outcomes were assigned significant weights and only 73.76% predictions of the best outcome had median errors below 100 km.
@@ 3.3.2 Ablation study metrics
Table 5: Performance results of the model trained and evaluated on the Great Britain + France + Canada dataset using the hybrid approach (Content + Context) and PMOP-type output of 5 prediction outcomes for the estimation of the tweet location.
Therefore, the ablation study on these specific countries effectively highlighted the model’s ability to handle ambiguity in language and demonstrated the influence of dataset composition on geolocation prediction accuracy. The ablation study results demonstrated that the model performed better on this limited dataset compared to the worldwide and American datasets. Notably, 96% of the predictions were within 161 km of the ground truth locations, which supports the suggestion [31] of training country-scale models.
@@ 4 Discussion and Conclusion
This work was aimed at the examination of the machine learning techniques for solving the short text geolocation prediction task with NLP techniques employing a scope of BERT-based neural networks. The proposed framework demonstrates simplicity in terms of the trained model application as a tool for geo-tagging textual big data of short corpora. While the process of finetuning the models involved some complexity in loss functions, it is a one-time step that requires initial time and computing resources to develop the methodology of training the most accurate models.
@@
For instance, setup requirements for the proposed model application are the following: access to the local or remote model at huggingface/geo-bert-multilingual; optionally dataset of text samples shorter than 512 tokens; and GPU or CPU node to launch the model.
@@
While the training routine requires access to the base model; dataset that contains labeled coordinates (lon, lat), text features for the arbitrary model input, and supplementary geo-tags of the locations; the usage of GPU is recommended due to the time and computing resources needed.
@@
One of the problems mentioned earlier was an alternative for the gazetteers—dictionaries of geographical indexes—used in the previous works containing accurate terms (location naming and LIWs) for geographical objects and their correct geographical coordinates. Considering the noisy user-generated data of the typical model inputs, the idea of feeding a consistent knowledge base of geographic objects into the model appeared practical and has been successfully implemented. The highest performance models were trained on multitask solving that offered an option of parallel learning with a total per-batch training loss as shown in Figure 7(d). The best strategy was split into key NON-GEO (always present ’text’ and ’user’) and minor GEO-ONLY (’place’ present only in geo-tagged tweets) text features. In practice, every feature had its individual linear regression layer wrapping the base model that differed in the number of prediction outcomes, but matched in the per-feature loss function type (Geospatial/Probabilistic).
The evaluation was performed using solely the key feature wrapper layer with no regard for the minor feature wrapper layer utilized only during the finetuning. This approach allowed for maintaining the format of the common model input and mapping the official place descriptions to the locations associated with the primary input texts. In terms of the described setup, the best performance results were achieved using the SOP-type minor loss and MOP-type key loss for 3–5 prediction outcomes.
Experiments have demonstrated that the geospatial loss of the minor feature GEO-ONLY shown in Figure 7(b) was declining much faster than the error distance of key feature NON-GEO shown in Figure 7(a) since the minor feature had less noisy and more persistent text data associated with the specific geolocations. Such behavior is expected in the case of the relatively high quality of metadata associated with the text which is generally a simpler problem than that of disambiguation within text documents in general [13].
The straightforward concatenation of the text content and metadata context of the ’place’ field would require undesirable randomization of all words/parts in the string before tokenization such that models would not learn to only pay attention to the optional part (’place’ metadata) of the input sequence. This statement is supported by the results in Table 4 which show that the model trained on ALL (union of ’text’, ’user’, and ’place’), as the key feature has underperformed the model trained on the data split to key and minor features. Moreover, the model trained solely on key NON-GEO feature with no regard to the place context demonstrated even higher spatial errors during the evaluation. Thus the proposed methodology of multitask learning has demonstrated a significant improvement in the accuracy of the predictions and should be utilized if it’s possible.
Another proposed novelty was a lower-bound limitation of the covariance parameter $\sigma_{\widehat{c}}$ in the GMM output of the probabilistic models. Since the probabilistic loss function computed the negative log-likelihood of the labeled location to fit in the predicted Gaussian distributions, it depended on the Probability Density Function (PDF). In the case, $\sigma_{\widehat{c}}$ was approaching its minimum at 0, PDF exceeded 1 driving the loss function to the negative values as shown in Figure 3. In terms of probabilistic models, the total loss depended on both spatial and probabilistic components, therefore the latter gained a bigger momentum. As a result, the model paid less attention to the squared Euclidean error distance described in Eq. (SED) which negatively affected the geospatial accuracy. The suggested solution limited PDF to [0, 1] interval by setting the minimum of $\sigma_{\widehat{c}}$ to $\frac{1}{2\pi}$ such that Probabilistic loss shown in Figure 7(c) stays in the positive domain. Since the covariance parameter measures model uncertainty about the predicted location point, in the best-case scenarios, the probabilistic loss would depend mainly on the distance error as described in Section C.
@@
Figure 7: Loss logged during the training of the proposed models with single and multiple (5) outcomes and Geospatial and Probabilistic loss; KF—Key Feature, MF—Minor Feature; model graph lines by color: PSOP, PMOP, GSOP, and GMOP where abbreviations refer to the configuration descriptions provided in Table 3.
@@
To address these privacy concerns, it is crucial for researchers, to anonymize or aggregate the data to minimize the risk of re-identification. The data utilized in this study has been protected with strict measures to prevent unauthorized access or breaches. Thus any collected Twitter dataset examples, containing metadata of user-profiles and geo-tags obtained from parsing the tweets, won’t be shared due to the potential privacy issues. By balancing the potential benefits of geolocation prediction with the protection of users’ privacy, it is possible to mitigate the privacy risks and promote responsible use of location data in Twitter research.
@@
In addition, there are several different world models that can be utilized for prediction output. While the Plate carrée projection is commonly used due to the Twitter geolocation stored in the WGS84 format, other projections such as the Mercator projection, Robinson projection, conic projection, and Winkel-Tripel projection should be considered as possible alternatives in future studies.
Besides the BERT variations, the proposed machine learning techniques have the potential to be applied to other base models. In the realm of NLP, various state-of-the-art models have been developed to enhance text analysis tasks, particularly text classification that provides an output representation for the examined regression task. However, this would require adapting the implementation of the wrapper layer to accommodate the shape of the pooler output vector specific to the chosen model. Currently, the approach is designed for linear regression logic, transforming the vector of size 768, which is common to all BERT-based models, into a predefined number of continuous numerical values. Furthermore, in order to ensure suitability for the downstream task of sentence classification, it is important that the base models possess strong contextual understanding and the ability to capture semantic relationships within sentences.
@@
Overall, location-based sentiment analysis is an important tool for understanding public opinion, social dynamics and patterns, helping decision-makers and researchers to make data-driven decisions, and support the work of various sectors, from marketing to governance. This study provides the NLP-based approach for the estimation of geolocation by processing short text corpora such as social media posts on Twitter. The proposed solution utilizes multitask learning (key and minor features), context data (user and place metadata), and probabilistic output (GMM) to achieve higher spatial accuracy on the tasks of the tweet and the user’s home location prediction. Thus contributing to the field of big data analysis with a flexible in geographical granularity setup of the custom BERT-based models.